an object thrown up with an instant velocity 40m/s from grounds. assume the energy is conservative. find the maximum height


1
Expert's answer
2021-01-06T13:55:20-0500

"E_k+Ep= const"

"mgh+\frac{mV^2}{2}= const"

"\text{where }E_k,E_p \text{kinetic and potential energy of the object}"

"g=9.8m\/s^2"

"\text{during the throw:}"

"h_0 = 0;V= V_0=40m\/s"

"E_k+Ep= mgh_0+\frac{mV_0^2}{2}=\frac{mV_0^2}{2}"

"\text{when the object reaches its maximum height:}"

"V= 0;E_k+Ep= mgh_{max}"

"\text{compare both parts}"

"mgh_{max}=\frac{mV_0^2}{2}"

"h_{max}=\frac{V_0^2}{2g}=\frac{40^2}{2*9.8}\approx16.66"

Answer:"h_{max}=16.66m"